A block of weight 7.0N rest on a level floor. The frictional force between the block and the floor is 1.0N. A horizontal force of 1.4N is used to pull the block for 4seconds. What is the velocity of the block after this time?


1
Expert's answer
2021-11-05T18:19:46-0400

The total force on the block is:


"F = 1.4N-1.0N = 0.4N"

where "-" sign is because the frictional force and pulling force are opposite.

Let the mass of the block be:


"m = \dfrac{7N}{9.8N/kg} = \dfrac{7}{9.8}kg"

where "9.8N\/kg" is the gravitational acceleration.

Then, according to the second Newton's law, the acceleration of the block is:


"a = \dfrac{F}{m}"

The velocity after time "t=4s" is given by the kinematic equation:


"v = at = \dfrac{Ft}{m}\\
v = \dfrac{0.4N\cdot 4s}{7/9.8kg} \approx 2.2 m/s"

Answer. 2.2 m/s.